Reports 1-1 of 1 Clear search Modify search
AdV-TCS (Ring Heater)
direnzo - 14:40 Monday 09 June 2025 (66940) Print this report
Comment to WE RH tuning (66928)

I'm posting an updated version of the plot showing the Omicron trigger rate from June 4 to today. In this version, I increased the averaging window from 1 hour to 3 hours to smooth out normal rate fluctuations and better highlight longer-term trends. I also added more color levels for the SNR to improve visualization of high SNR glitches.

In Figure 1, the top panel shows the usual time-frequency glitchgram. The bottom panel displays the glitch rate per minute, with different colored curves corresponding to different SNR thresholds. The DQ bar indicates segments in DQ Studies mode (light green) and Low Noise 3 (dark green).

A decrease in glitch rate is visible starting from June 7, likely associated with the WE RH power increase from 7W to 7.3W: #66928 . There also seems to be a gradual downward trend in the rate between June 7 and today.
These effects are more evident in the upper curves, hence corresponding to lower SNRs (SNR < 200).

For comparison, I also include a version of the rate plot averaged over one hour (Figure 2), as for the Omicron plot previously posted by Nicolas.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×